# A NIRCam-dark galaxy detected with the MIRI/F1000W filter in the MIDIS/JADES Hubble Ultra Deep Field

### v1

@@ 3.2 Multi-wavelength validation and observational properties of the NIRCam-dark object Cerberus
For the MUSE data, we made use of the recent MXDF observations that cover the XDF area with observations in WFM mode for a total of more than 140 hours (see Bacon et al., 2023, for more details). We downloaded the publicly available fully reduced MUSE data, as well as a catalogue of detected sources and corresponding spectroscopic redshifts111The MUSE cube and catalogue of sources can be found at the following link: https://amused.univ-lyon1.fr/project/UDF/HUDF/.. We extracted the MUSE 1D spectrum at the coordinates of our target within circular apertures of different sizes (04 and 08 radius) but found no significant spectral feature at the position of the NIRCam-dark candidate. This is also confirmed by the publicly available catalogue of detected sources by Bacon et al. (2023).
We also looked for emission lines in the MUSE data in the neighborhood of our source. We detected extended Ly emission from a nearby source lying at and celestial coordinates = 03:32:37.95 (hours), = 27:47:11.14 (degrees). This source is reported by Bacon et al. (2023) with ID53. Interestingly, according to Bacon et al. (2023), there is Ly emission associated with another source (ID7666), that we do not detect in the MXDF datacube, at with celestial coordinates = 03:32:38.08 (hours), = 27:47:12.71 (degrees). Both objects are at an angular separation of less than 1″ from our target.
Searching for a detection at (sub-)millimiter wavelengths, we made use of the deep ALMA dust continuum imaging from the ASPECS program at 1.2 mm and 3.0 mm that covers the HUDF. By carefully inspecting these maps, the source is not detected at either wavelength. This finding implies upper limits on the flux density at 1.2 mm and 3.0 mm of 27.9 Jy and 4.2 Jy respectively, assuming it is effectively an unresolved point-like source. Figure 4 presents the images of Cerberus in the ALMA 1.2 and 3 mm bands.
@@ 4.1 Multi-wavelength Photometry
A spectral energy distribution (SED) for Cerberus was built by measuring photometry in all available NIRCam (including both individual bands and stacks) and MIRI bands. Given the faint and small nature of our source, we measured photometry in several circular apertures with radii ranging from 0.1″ to 0.6″. We measured the background in a small square region around Cerberus (3″ on the side) and calculated the noise using non-adjacent pixels (5 pixels apart) to take into account noise correlation, as explained in Pérez-González et al. (2023b).
All NIRCam SW measurements provided values lower than 2. The 5 depths for point-like sources of the 6 SW filters range from 29.6 mag for F090W to 30.6 mag for F200W, measured in each case with an aperture enclosing 65% of the encircled energy ( radius).
The measurements in the NIRCam LW channels presented for all apertures, only the F277W and F460M filters approached that limit. For these bands, we obtained mag and mag (after removing the cosmic-ray artifact mention in Section 2), respectively, measured in a 02 radius aperture, and accounting for an 0.3 mag aperture correction. Lower significances but near the limit were obtained in F335M, F410M, F444W, and F480M, mag, mag, mag, and mag, respectively.
@@
Given our magnitude measurement in the F1000W band, mag, obtained with photometry in an aperture of radius 03, which maximizes the , we infer a very red color, 0.4 mag.
@@
These measurements and the stamps provided in Figure 2 show that Cerberus is barely detected in any individual NIRCam band, consistent with being around or below the 5 detections limits, quoted to be around 30.5–31.0 mag for the JADES DR2 data (Eisenstein et al., 2023a). We remark that the F460M filter, which provides the brightest magnitude, presents a group of 4 pixels to the North of the position of our galaxy, which we confirmed to be an artifact of the reduction, and it is not seen in any other band (even the ones with deeper limiting magnitudes).
@@
Figure 2 also shows stacked NIRCam data for Cerberus. Visually, it seems clear that there is signal at that sky position. We measured photometry in the same apertures used for the individual bands, assigning average aperture corrections for the filters included in each stack. The 5 LW stacks with the highest number of bands all provide fluxes (the average being ), ranging from 30.3 to 30.8 mag. This means that the (low-S/N) SED is characterized by a red color extending through all the stacks from 2 to 4 m and beyond, up to 0.3 mag for the first (reddest) stack (including F480MF444W), which reaches m. For the SW stacks, no significant signal is recovered in any of them. We measure a upper limit of 31.0 mag for the stack adding all SW data. This implies an 0.7 mag color jump from the NIRCam SW to the LW ranges.
@@
The SED of Cerberus shown in Figure 5 is, in fact, very similar to the little red dots (LRDs) recently presented and characterized in a number of papers (e.g., Labbé et al. 2023a, b, Matthee et al. 2023, Barro et al. 2023, Williams et al. 2023a, Pérez-González et al. 2024, Kokorev et al. 2024), but shifted to longer wavelengths. The mag general selection criterion of LRDs, and, more specifically, the 3 mag color of the example source 203749 in Pérez-González et al. (2024), compares well, if we multiply the observed wavelengths by a factor of 2, with the mag (or, using the F560W upper limit, mag). The flat F150WF200W color distinctive of LRDs (the largest value allowed by the selection is typically 0.5 mag), i.e., a change in slope at around 0.3-0.4 m in the rest-frame for the typical redshift of LRDs , compares well with the color between 3 and 4 m for Cerberus, just multiplying again the observed wavelengths by a factor of 2. That rest-frame UV color is indeed consistent with the reddest LRDs in Pérez-González et al. (2024) or the LRD of which a NIRSpec spectrum was presented in Killi et al. (2023). The analogy can also be established in terms of the apparent magnitude: the typical magnitude of LRDs in the blue part of their SED lies around 27-28 mag, Cerberus being times dimmer in NIRCam LW bands. We note that the difference in luminosity distance and cosmological dimming between , the median redshift of LRDs (Pérez-González et al., 2024) and (a factor of 2 larger, see also Section 4.2.3) will translate to a flux ratio around an order of magnitude for the same type of object.
@@ 4.2 Analysis of the SED of Cerberus
In the following subsections, we discuss possible interpretations of the SED of Cerberus. The NIRCam and MIRI source (Figure 3) were observed more than one year apart. However, given the spatial coincidence of the object in the NIRCam and MIRI observations (centroids separated by 0016), we proceed by assuming that they are the same object and not a chance alignment of two different objects.
@@ 4.2.2 Galactic origin
As shown by Langeroodi & Hjorth (2023a) and Hainline et al. (2023) for LRDs, red sources detected by JWST in cosmological fields such as GOODS-S can be brown dwarfs. Following the method in Langeroodi & Hjorth (2023a), we fit the photometry of Cerberus with brown dwarf atmosphere templates. We include the Sonora cloud-free models (Karalidi et al., 2021; Marley et al., 2021) as well as the ATMO2020++ cloud-free adiabat-adjusted T and Y dwarf models (Leggett et al., 2021; Meisner et al., 2023, see also Suárez & Metchev 2022). In particular the ATMO2020++ models were included to ensure coverage of extremely cool brown dwarfs with 250 K 500K. Compared to warmer brown dwarfs, these models more closely resemble the SED shape of Cerberus, which is relatively faint at near-infrared wavelengths while getting significantly brighter at 10 m.
The best-fit brown dwarf template is shown in the top-left panel of Figure 6. We infer a best-fit effective temperature ( [K]) of , surface gravity ( [cm s]) of , and distance of pc. Its effective temperature would most likely classify Cerberus as a Y-type dwarf (see Figure 12 in Langeroodi & Hjorth, 2023a). Compared to the UltracoolSheet compilation of known brown dwarfs with parallax distances (Best et al., 2020), its inferred distance (pc) would make Cerberus one of the farthest (top 20) brown dwarfs discovered to date (see Figure 11 in Langeroodi & Hjorth, 2023a). This distance at the high Galactic latitude of Cerberus would place it at the edge of the thin disk (see Figure 6 in Hainline et al., 2023). Recently, three T-type dwarfs were discovered and spectroscopically confirmed with NIRSpec prism (Langeroodi & Hjorth, 2023a; Burgasser et al., 2023) at estimated distances between 0.7 and 4.8 kpc, all most likely outside the thin disk; among them A2744-BD3 is a late T-type at 755 pc.
@@
In addition, both the MIRI and stacked NIRCam images (see Section 3.2) shows the source to be slightly extended and asymmetrical as compared to the PSF. The two instruments, NIRCam and MIRI, indicate a morphology for the source that makes it unlikely that it is a point source, thus disfavoring the single brown dwarf scenario. The slight extension of order alone would however not argue against a brown dwarf binary. Even at only 100 pc distance, a semi-major axis of for a binary of two identical brown dwarfs/planets with masses of M each would result in orbit periods of order 100–1000 years, i.e., the elongated image would not vary between the NIRCam and the MIRI observations.
4.2.3 Extragalactic origin
The SED of Cerberus was fitted to estimate a photometric redshift and obtain its physical properties in a 2-step procedure. We used several codes for this purpose, separating the two tasks to analyze facets of each method and understand the (possibly systematic) uncertainties in detail. We used the fluxes plotted in Figure 5 as well as the ALMA upper limits quoted in Section 2, shown to be important for high redshift sources (even for determining their redshift, see Meyer et al. 2024), and particularly for an ERO such as Cerberus. The results are shown in Figures 6 and 7.
We first used the eazy code (Brammer et al., 2008) to fit the data in 2 ways: one using actual (low ) flux measurements for all filters, and another replacing all measurements by upper limits that templates were not allowed to surpass (achieved with a modified version of the code). We employed in the fits all v1.3 templates, which include a dusty galaxy with a a high-equivalent width emission-line spectrum, as well as new models for little red dots and high-redshift AGNtorus recently added based on JWST data (Killi et al., 2023). We did not impose any prior and worked with minimum photometric redshift estimates in the range .
The synthesizer-AGN code assumes that the SED can be modeled with a composite stellar population (Pérez-González et al., 2003, 2008) and AGN emission coming from the accretion disk and the dust torus (Pérez-González et al., 2024). The stellar emission includes a young and a more evolved star formation event, each one described by a delayed exponential function with timescales between 1 Myr and 1 Gyr, and with ages from 1 Myr up to the age of the Universe at the redshift of the source. The attenuations of the emission from each stellar population are independent and described by a Calzetti et al. (2000) law, with A values ranging from 0 to 10 mag for each population. The stellar emission is described by the Bruzual & Charlot (2003) models, assuming a Chabrier (2003) IMF with stellar mass limits between 0.1 and 100 , and the nebular emission is also considered (Pérez-González et al. 2003). The AGN emission is modeled with a QSO average spectrum (Vanden Berk et al., 2001; Glikman et al., 2006). The dust emission from the AGN is modeled with the self-consistent templates of AGN tori presented in Siebenmorgen et al. (2015), and dust emission linked to star formation is added using the models from Draine & Li (2007).
CIGALE (Boquien et al., 2019) SED-fitting analysis assume an SFH modelled using a constant star-formation rate with ages ranging from 1 to 100 Myr. We adopted the stellar populations models from Bruzual & Charlot (2003) with solar metallicity, and the Chabrier (2003) IMF. We included nebular continuum and emission lines using solar metallicity, electron density of 100 cm, and ionised parameter equal to =2. The dust attenuation and far-IR emission uses the Calzetti Law (Calzetti et al., 2000) and the Draine et al. (2014) models, respectively. AGN emission is added using the Fritz et al. (2006) models following the initial parameters suggested by Ciesla et al. (2015).
The results of our photometric redshift analysis are given in Figure 5. All codes provide 2 main solutions and a number of tertiary redshift values. We mark the 2 main solutions at and for each code in the right panel of the figure. The obtained solutions are: (provided by prospector and LePHARE++), (very prominent solution for by eazy, also detected weakly by prospector and LePHARE++), (obtained by all codes), (obtained by LePHARE++), and (obtained by bagpipes and LePHARE++). We note that the solution corresponds to a PAH line entering the F1000W passband, the solution would imply Pa covered by the filter, for , H would contribute to the MIRI-F1000W flux, and means that the F1000W emission could be enhanced by H and/or [O iii]4960,5008 emission.
After identifying all possible solutions, we fixed the redshifts to the peaks showing the highest probabilities, namely , , and , and ran all codes to understand the implications on the nature of Cerberus. We analyze in detail the different possibilities to explain the SED of Cerberus in the following subsections. We provide best-fitting models in Figures 6 (focused on the spectral range typically dominated by stellar emission) and 7 (extending the SED plots to the spectral region that might be dominated by dust emission). The results are summarized in Table 1.
High- solution:
All codes provide a photometric redshift solution around –. As mentioned before, this solution would correspond to H entering the F1000W filter, which would leave H and [O iii]4960,5008 within the F770W filter.
In the absence of dust, the F444WF1000W color is very difficult to reproduce with stars alone, even for the youngest and lowest metallicity models. Taking into account the width of the F1000W filter, even a H line with a prominent Å (the limit for Starburst 99 models, Leitherer et al. 2002, using the lowest metallicities, the youngest ages, and setting all relevant parameters to their extreme values, e.g., ionizing photon escape fraction) would produce a relatively small flux density enhancement in the F1000W filter (around 1 mag). The only way to fit the full SED is including some red component, either a dusty starburst or an obscured AGN.
Indeed, the best fits for this solution are obtained with two different models. The first possibility is a relatively dusty ( mag for all codes except prospector, but we remark that this software internally increases the attenuation for younger stars), massive ( M), and intermediate age (50–100 Myr mass-weighted age) starburst. Those numbers would translate to SFRs around 0.1–30 M yr.
The second possibility is a model where a significant fraction of the F1000W flux comes from the emission from a dust torus, and there is some contribution to the SED coming from star formation, increasing towards the ultraviolet range. This solution allows a significantly smaller stellar mass, M, for synthesizer-AGN (with large uncertainties due to the overshining effect of the AGN) in a young (1 Myr age, 20 M yr, mag) star-forming galaxy. The other codes including an AGN, bagpipes and cigale, provide similar results, although the stellar attenuation is significantly higher, and the fits present problems in reproducing the ALMA upper limits (see Figure 7).
The obscured AGN fit to Cerberus SED resembles the results for the reddest LRDs at , but within a 100 times less massive galaxy and at a 500(250) Myr earlier epoch compared to the average (maximum) redshift of LRDs (Pérez-González et al., 2024; Greene et al., 2023; Killi et al., 2023).
The derived bolometric luminosity of Cerberus for the synthesizer-AGN solution is erg s. Given the high redshift, the very red NIRCam-to-MIRI color, and the ALMA constraints, the torus models favored by this SED-fitting code peak at very short wavelengths (20 m rest-frame, 300 m observed), as is shown in Figure 7. Translating those extreme conditions to a near- or even above-Eddington limit accretion, we could estimate the presence of a 10 M SMBH using the results in Woo & Urry (2002). We remark that these are very uncertain calculations because they are based on the F444WF1000W measured color probing the bluest emission of the possibly present torus and the ALMA upper limits which probe the Rayleigh-Jeans tail of the torus emission. We should also consider that variations of the amount of stellar and nebular emission strongly affect the bolometric luminosity calculation. Taking all the numbers above, and considering the mentioned caveats, the stellar to SMBH mass ratio would be around 0.2, consistent with the values discussed in Pacucci et al. (2023, see also ); Pacucci & Loeb (2024, see also ) for high redshift AGN discovered with JWST.
Intermediate- solution:
A solution at is the main one provided by eazy, which could correspond to Pa entering the F1000W passband (for ). However, the best eazy template corresponds to a dusty galaxy (again, the NIRCam-MIRI color is too large for any other model). The dusty nature is confirmed by the more detailed SED modeling performed with synthesizer-AGN after fixing the redshift, which reproduces the SED with a 200 M yr, mag, 5 Myr old starburst in a M galaxy. bagpipes obtains very different solutions for the AGN and SF versions: a starburst (10 Myr old galaxy) M galaxy with large amounts of dust, mag, or an obscured AGN in a more massive post-starburst galaxy, M, with older age and larger dust content. Similar properties are obtained with other codes, pointing to a dusty galaxy with a stellar mass one order of magnitude smaller than other analog sources discovered and/or analyzed with JWST (consistent with their brightest NIRCam magnitudes, see, e.g., Barrufet et al. 2023, Pérez-González et al. 2023a, or Williams et al. 2023a).
Within this redshift range, we should consider the possibility of Cerberus being an exotic type of faint, heavily obscured AGN whose SED differs substantially from the bright quasar templates included in our fitting codes (although it could be recovered with a QSOtorus composite by synthesizer-AGN). In particular, Hamann et al. (2017) report ERQs at whose flux decrement between 12 m and 1 m is consistent with Cerberus. ERQs are characterized by faint and intrinsically red optical continua and abnormally strong emission lines and infrared emission; their SEDs are not part of standard AGN fitting libraries due to extreme rarity. However, all reported ERQs are brighter than Cerberus by 14 magnitudes. The AGN processes leading to ERQs are unknown, as is whether they can be sustained at much lower bolometric luminosities. Therefore, this possibility cannot be assessed quantitatively.
We note here that Cerberus is located very near (″) 2 Lyman Alpha Emitters at , as presented in Bacon et al. (2023): ID7666 at , and ID53 at . Besides, considering all the sources found in the MXDF between reported in Bacon et al. (2023) with reliable spectroscopic redshifts, our target lies in an overdense region of Lyman Alpha Emitters within a radius of 15 arcsec (i.e. about 100 pkpc). Within the 15 arcsec radius aperture, we find 17 galaxies between , 10 galaxies at , 5 objects at and 20 galaxies at . This points to a somewhat similar case as the galaxy claimed to lie at in the CEERS field (Donnan et al. 2023; Finkelstein et al. 2023; although other photometric redshift estimations provided , Pérez-González et al. 2023a) but being part of a large scale structure (Naidu et al., 2022, see also Zavala et al. 2023 for a discussion about dusty galaxies with and photometric redshift solutions) and finally confirmed to be at (Arrabal Haro et al., 2023). Several cases like this one have been reported recently (Caputi et al., 2021; Kokorev et al., 2023).
Low- solution:
For the last photometric redshift solution, , in this case provided by prospector and LePHARE++, we fixed the value for detailed SED modelling at . The F1000W flux here is dominated by the 7 m PAH feature. This solution then implies a highly attenuated ( mag) dwarf galaxy for all codes with a stellar mass M. Alternatively, bagpipes-AGN reproduces the SED with an obscured AGN in a galaxy with a similar (small) stellar mass. The very small stellar mass, relatively high dust attenuation, and strong PAH emission is an uncommon combination, at odds with the mass-metallicity relationship (e.g., Tremonti et al., 2004) and the weakening of PAH features at low metallicities (e.g., Engelbracht et al., 2005).
4.2.4 Transient nature
The bright nature of Cerberus hints at a possible transient nature of the source. In fact, analyzing the 11 individual MIRI exposures (taken within a 10 hours window), and partial mosaics of the full dataset, we found that excluding 2 exposures from the stack implied that the source becomes 0.3–0.4 mag dimmer. The origin of this difference was linked to 2 contiguous exposures, numbers 7 and 8. This is a small difference (within 2 of the quoted F1000W magnitude), but the fact that the 2 exposures are contiguous might be linked to the source having a transient nature. Alternatively, it might have originated from some transient effect on the detector, for example, cosmic ray showers (although no hit was detected by the pipeline). If the origin is astrophysical, it could be explained in terms of variability of a rotating asteroid (Sanchez et al., 2017, although we do not detect any proper motion within a year, see Section 4.2.1), a brown dwarf (e.g., Artigau et al., 2009), or an AGN (Sheng et al., 2017).

### v2

@@ 3.2.2 MUSE measurements
We extracted the MUSE 1D spectrum at the coordinates of our target within circular apertures of different sizes (04 and 08 radius) but found no significant spectral feature at the position of the NIRCam-dark candidate. This is also confirmed by the publicly available catalog of detected sources by Bacon et al. (2023).
3.2.3 ALMA measurements
The source is not detected at 1.2 mm and 3.0 mm. This finding implies upper limits on the flux density at 1.2 mm and 3.0 mm of 47 Jy and 7 Jy respectively, assuming it is effectively an unresolved point-like source.
@@ 4.1 Multi-wavelength Photometry
A spectral energy distribution (SED) for Cerberus was built by measuring photometry in all available NIRCam (including both individual bands and stacks) and MIRI data, forcing measurements at the position determined for Cerberus with the MIRI data. Given the faint and small nature of our source, we measured forced photometry in several circular apertures with radii ranging from 0.1″ to 0.6″. We measured the background in a small square region around Cerberus (3″ on the side) and calculated the noise using non-adjacent pixels (5 pixels apart) to take into account noise correlation, as explained in Pérez-González et al. (2023b), obtaining similar depth results (within 0.2 mag) to those reported in Eisenstein et al. (2023a) for the NIRCam bands. Photometric measurements are provided in Table 1.
None of the measurements in individual NIRCam bands have a significance, which we use as our limit to define a detection in each filter.
@@
Given our magnitude measurement in the F1000W band, mag, obtained with photometry in an aperture of radius 03, which maximizes the , and the measurements in individual bands and in stacked data for the reddest LW NIRCam channels, we infer a very red color, 0.4 mag. Using the reddest disjoint stack with an actual detection (i.e., ), we obtain a color mag.
@@
These measurements indicate that Cerberus is not robustly detected in any individual NIRCam filter, but consistent with being slightly below the 5 detections limits in some bands, quoted to be around 30–31 mag for the JADES DR2 data (Eisenstein et al., 2023a). We remark that the F460M filter, which provides the brightest magnitude, presents a group of 4 pixels to the North of the position of our galaxy, which we confirmed to be an artifact of the reduction, and it is not seen in any other band (even the ones with deeper limiting magnitudes). We excluded this filter from the rest of analysis.
@@
We measured photometry in the stacked data in the same apertures used for the individual bands, assigning average aperture corrections for the filters included in each stack. The 5 LW stacks with the highest number of bands all provide fluxes (the average being ), ranging from 29.9 to 30.8 mag. This means that the (low-S/N) SED is characterized by a red color extending through all the stacks from 2 to 4 m and beyond, up to 0.2 mag for the first (reddest) stack (including F480MF444W), which reaches m. For the SW stacks, no significant signal is recovered in any of them. We measure a upper limit of mag for the stack adding all SW data. The comparison between SW and LW stacks implies a 0.6 mag color jump from the to the m ranges.
@@
The SED of Cerberus is, in fact, very similar to the little red dots (LRDs) recently presented and characterized in a number of papers (e.g., Labbé et al. 2023a, b, Matthee et al. 2023, Barro et al. 2023, Williams et al. 2023a, Pérez-González et al. 2024, Kokorev et al. 2024), but shifted to longer wavelengths. Based on detections of Cerberus in stacked images, we can infer that the mag general selection criterion of LRDs, and, more specifically, the 3 mag color of the example source 203749 in Pérez-González et al. (2024), compares well, if we multiply the observed wavelengths by a factor of 2, with the mag, or, using the F560W upper limit, mag. The flat F150WF200W color distinctive of LRDs (the largest value allowed by the selection is typically 0.5 mag), i.e., a change in slope at around 0.3-0.4 m in the rest-frame for the typical redshift of LRDs , compares well with the color between 3 and 4 m for Cerberus, just multiplying again the observed wavelengths by a factor of 2. We remark that this comparison is based on S/N5 measurements in stacked data. That rest-frame UV color is indeed consistent with the reddest LRDs in Pérez-González et al. (2024) or the LRD of which a NIRSpec spectrum was presented in Killi et al. (2023). The analogy can also be established in terms of the apparent magnitude: the typical magnitude of LRDs in the blue part of their SED lies around 27-28 mag, Cerberus being times dimmer in NIRCam LW bands. We note that the difference in luminosity distance between , the median redshift of LRDs (Pérez-González et al., 2024) and (a factor of 2 larger, see also Section 4.2.1) will translate to a flux ratio between bands covering the same rest-frame wavelength around an order of magnitude for the same type of object.
@@ 4.2 Analysis of the SED of Cerberus
In the following subsections, we discuss possible interpretations of the SED of Cerberus. The NIRCam and MIRI datasets were taken more than one year apart. However, given the spatial coincidence of the object in the NIRCam and MIRI observations (centroids separated by 0016), we proceed by assuming that they are the same object and not a chance alignment of two different objects.
4.2.1 Extragalactic origin
The SED of Cerberus was fitted to estimate a photometric redshift and obtain its physical properties in a 2-step procedure. We used several codes for this purpose, separating the two tasks to analyze facets of each method and understand the (possibly systematic) uncertainties in detail. We used the fluxes described above as well as the ALMA upper limits quoted in Section 2, shown to be important for high redshift sources (even for determining their redshift, see Meyer et al. 2024), and particularly for an ERO such as Cerberus. We provide details about the fitting methods in Appendix A.
The results of our SED modeling analysis are given in Figures 3 (photometric redshift probability distributions) and 4 (SED models).
We remark that the SED of Cerberus is very noisy, given that it is an ultra-faint source at the very depth edge of the JWST surveys. Therefore, the physical properties of the source are highly uncertain, starting from its redshift (see several peaks in Figure 3).
However, leaving the redshift apart, there are two physical properties for which our results are similar among all SED modelling codes, mainly because they are linked to the 2 main observational characteristics of Cerberus, i.e., its faintness (translating to relatively low luminosity) and red color. The very red color of our NIRCam-dark object can only be reproduced with significant amounts of dust, either linked to PAHs if the redshift is low, , or warm/hot dust if the redshift is around or . For the lowest redshift, all codes provide very low stellar masses, around M⊙. The very small stellar mass, relatively high dust attenuation, and strong PAH emission is an uncommon combination, at odds with the mass-metallicity relationship (e.g., Tremonti et al., 2004) and the weakening of PAH features at low metallicities (e.g., Engelbracht et al., 2005). For the higher redshift solutions, and , all codes provide masses around M⊙, and they all need large warm dust contents, which can be explained with a very compact starburst or an AGN torus (similar to those detected at high redshift by JWST, see Maiolino et al. 2023, as well as Pacucci et al. 2023; Pacucci & Loeb 2024, and also Hamann et al. 2017).
@@ 4.2.3 Galactic origin
As shown by Langeroodi & Hjorth (2023a), Burgasser et al. (2024), and Hainline et al. (2023) for LRDs, red sources detected by JWST in cosmological fields such as GOODS-S or A2744 can be brown dwarfs. Following the method in Langeroodi & Hjorth (2023a), we fit the photometry of Cerberus with brown dwarf atmosphere templates. We include the Sonora cloud-free models (Karalidi et al., 2021; Marley et al., 2021) as well as the ATMO2020++ cloud-free adiabat-adjusted T and Y dwarf models (Leggett et al., 2021; Meisner et al., 2023, see also Suárez & Metchev 2022). In particular the ATMO2020++ models were included to ensure coverage of extremely cool brown dwarfs with 250 K 500K. Compared to warmer brown dwarfs, these models more closely resemble the SED shape of Cerberus, which is relatively faint at near-infrared wavelengths while getting significantly brighter at 10 m.
The best-fit brown dwarf template is shown in the top-left panel of Figure 4. We infer a best-fit effective temperature ( [K]) of , surface gravity ( [cm s-2]) of , and distance of pc. Its effective temperature would most likely classify Cerberus as a Y-type dwarf (see Figure 12 in Langeroodi & Hjorth, 2023a). Compared to the UltracoolSheet compilation of known brown dwarfs with parallax distances (Best et al., 2020), its inferred distance (pc) would make Cerberus one of the farthest (top 20) brown dwarfs discovered to date (see Figure 11 in Langeroodi & Hjorth, 2023a). This distance at the high Galactic latitude of Cerberus would place it at the edge of the thin disk (see Figure 6 in Hainline et al., 2023). Recently, three T-type dwarfs were discovered and spectroscopically confirmed with NIRSpec prism (Langeroodi & Hjorth, 2023a; Burgasser et al., 2023) at estimated distances between 0.7 and 4.8 kpc, all most likely outside the thin disk; among them A2744-BD3 is a late T-type at 755 pc.
@@
In addition, both the MIRI and stacked NIRCam images (see Section 3.2) shows the source to be slightly extended and asymmetrical as compared to the PSF. The two instruments, NIRCam and MIRI, indicate a morphology for the source that makes it unlikely that it is a point source, thus disfavoring the single brown dwarf scenario.
The slight extension of order alone would however not argue against a brown dwarf binary. Even at only 100 pc distance, a semi-major axis of for a binary of two identical brown dwarfs/planets with masses of M⊙ each would result in orbit periods of order 100–1000 years, i.e., the elongated image would not vary between the NIRCam and the MIRI observations.